Given 2p-x=50 and p=8000/(x+370), how can you find x?
Answer by stanbon(75887)

Given 2p-x=50 and p=8000/(x+370), how can you find x?
---------------------------------------
Substitute the "x expression" in place of "p".
-------
2[8000/(x+370)] - x = 50
16000/(x+370) - x = 50
16000 - x(x+370) = 50(x+370)
16000 -x^2+370x = 50x+ 18500
x^2 -320x + 2500 = 0
x = [320 +-sqrt(320^2-4*2500)]/2
x = [320 +- sqrt(92400)]/2
x = [320 +- sqrt(400*231)/2
x = [320 + - 20sqrt(231)]/2
x = [160 +- 10sqrt(231)]
x = 160 + 10sqrt(231) or x = 160 - 10sqrt(231)
Cheers,
Stan H.